OpenStudy (anonymous):

Will fan and give medal to who ever answers the following question: Margie is responsible for buying a week's supply of food and medication for the gogs and cats at a local shelter. The food and medication for each dog costs twice as much as those supplies for a cat. She needs to feed 164 cats and 24 dogs. Her budget is $4240. How much can Margie spend on each dog for food and medication?

OpenStudy (anonymous):

164c + 24(2c) = 4240 164c + 48c = 4240 212c = 4240 c = 20 She can spend $20 on each cat. d = 2c d = 2(20) = 40 She can spend $40 on each dog